these two triangle are similar solve for x

|\ /|
| \ 38 / |x+7
3x+3 | \ / |
^ \ / ^
| \ / |
| \ / |
| \ / |
|_______\ /________|
19
the two triangle touch each other

Answer by Alan3354(69443)

these two triangles are similar
solve for x
-----------
The triangles are similar.
The 38 side is twice the 19 side --> 3x+3 = 2*(x+7)
3x+3 = 2x + 14
x = 11